OpenStudy (heyitszia1):

FAN AND MEDAL!! The function f(x) = −x² + 48x − 432 models the profit, in dollars, a game developer makes for selling a new game, where x is the number of games sold and f(x) is the amount of profit. Part A: Determine the vertex. What does this calculation mean in the context of the problem? (5 points) Part B: Determine the x-intercepts. What do these values mean in the context of the problem? (5 points)

OpenStudy (heyitszia1):

I know that the vertex is (24,144), but I need help finding the x-intercepts.

OpenStudy (heyitszia1):

#math&ing001

OpenStudy (heyitszia1):

@Whitemonsterbunny17

OpenStudy (heyitszia1):

@Aveline

OpenStudy (heyitszia1):

I need help with part B

OpenStudy (aveline):

−x² + 48x − 432 To find the vertex, you use -b/2a a=-1 b=48 c=-432 So, the x value of the vertex is equal to: -48/-2 (24,144)

OpenStudy (aveline):

Part B: Determine the x-intercepts. What do these values mean in the context of the problem? (5 points) Let's simplify the equation: \[x ^{2}-48x+432=0\] \[(x-12)(x-36)=0\]

OpenStudy (aveline):

This means that the x-intercepts are at 12 and 36

OpenStudy (aveline):

I divided by -1 to make -x^2 positive

OpenStudy (heyitszia1):

Did you factor the equation to get (x - 12) and (x - 36)?

OpenStudy (aveline):

yes

OpenStudy (heyitszia1):

Thanks :)